OpenStudy (anonymous):

Write an equation of a line that passes through the points (-2,5) and (1,2). Help?

OpenStudy (anonymous):

y1-y2/x1-x2 this will find slope once you have ur slope the asnwer is y=(slope)x+-2

OpenStudy (anonymous):

Can you show your work please that would be helpful

OpenStudy (anonymous):

It's (y2-y1)/(x2-x1). Doing this will give you (2-5)/(1+2) = -1. then choose any set of points and use the point slope equation. for example, y-5=-1(x+2). That simplifies to be y-5=-x -2, ultimately becoming y=-x+3